# Distributed Multisensor ISAC

### v1

@@ VI Synchronization, Data Fusion, and Tracking in Distributed MS ICAS Networks
Advanced MS ISAC is highly cooperative. This begins with the bistatic sensing architecture, which is quite close to the basic radio access scheme of mobile communication. It already has all the advantages of bistatic radar w.r.t. target detection and identification. We refer to this first level as (i) Sensor Level Cooperation (SLC), which refers to processes that can be performed by each sensor pair independently of each other. Although it does not involve the exchange of processed data with other sensor nodes, it is already cooperative, e.g. when the sensor receives measurement signals from several illuminating radio nodes, e.g. in the MS ISAC up-link mode. The next level of cooperation is achieved when multiple sensing links share processed data to estimate and track the 3D dynamic target state vector of one or more targets. This level already provides important benefits of MS ISAC, which include spatially smooth performance distribution by multisensor data fusion, multilink access coordination, and local radio resource allocation. According to our initial discussion on MS ISAC architectures, the ISAC nodes may belong to a distributed infrastructure as an MS sensing cluster site, include mobile UEs (in DL or UL), or belong to an independent, self-contained cloud of meshed radio nodes. We refer to this level as (ii) Intra Site (or Intra Cloud) Cooperation (IntraSC). The third level of cooperation is called (iii) Intersite (or Inter Cloud) Cooperation (InterSC). It fuses the estimated target state vectors from different sensing sites or clouds and can combine or fuse ISAC data with other (including non ISAC) information to produce a comprehensive situation report. In the sequel we will at first discuss some synchronization issues and then shortly refer to distributed data fusion and tracking in MS ISAC.
VI-A Differential Coherent Operation in MS ISAC at Sensor Level
@@
At sensor level, parameters are estimated that are closely related to the dynamic target state vector, such as time delay (rsp. range), Doppler frequency shift (rsp. relative speed) or directions (DoD, DoA). While the latter in general requires more complicated RF interfaces (because of the antenna arrays), the former two require synchronization with respect to sampling and Local Oscillator (LO) frequencies offsets of the bistatic radio nodes [ref90]. Array processing is easier in this respect as only local (or platform related) coherent processing is necessary. Therefore, distributed MS-MIMO ISAC is often claimed to have disadvantages due to a lack of synchronization between sensor nodes. On closer inspection, however, many solutions can be found if we avoid excessive and unnecessary requirements. As already explained above, MS-MIMO ISAC requires special efforts to establish sensor access coordination and synchronization. However, since the Tx and Rx are connected as a communication transceiver pair in a multiuser or mesh communication network, the receiving radio nodes are basically synchronized to the transmitter by using the Primary and Secondary Synchronization Signals (PSS, SSS) [ref91, ref92]. Therefore multisensor operation, OFDM frame coordination and carrier frequency adjustment is already accomplished according to the requirements for communication. This ensures orthogonal, mutual interference free medium access and synchronous FFT processing which is a fundamental advantage for low variance system identification [ref30]. This is further supported by the design the OFDM symbol, which is not chosen too long in order to avoid time selectivity and supplemented by a cyclic prefix. For further synchronization update one to four Demodulation and Phase Tracking Reference Signals (PT-RS und DM-RS) can be allocated per slot within the Physical Downlink / Uplink Shared Channel (PDSCH/PUSCH) [ref91, ref92].
@@
Although this timing and clock synchronization schemes are very flexible and powerful for communication even in situations with high Doppler shift, they seem not to be enough for ISAC on the sensor level cooperation. ToF and Doppler estimation call for sub-ns delay and precise carrier phase synchronization (the latter becoming more demanding the higher the frequency). Fortunately, these problems can be mitigated by carrying out differential measurements. A striking historical example is given by the well-known passive radar concept, which uses a “transmitter of opportunity” as an illuminator and a “passive” (receiving only) sensor. PCL does not presume any explicit signaling scheme for synchronization between Tx and Rx. The bistatic eToF is estimated from the received communication signal, which is correlated against the LOS-reference. Classical passive radar receivers use two synchronous channels, an omnidirectional surveillance channel and a directive channel receiving the clean LOS reference, which must be free of multipath. In case of the more advanced CPCL, the eToF and eDoppler can be actually calculated from the same CIR by the delay and frequency difference between LOS and BTP, provided that the correlation reference is correctly reconstructed and LOS and target peaks are clearly identified. This eliminates the remaining Tx and Rx clock and LO offset and even avoids drift and deviations between parallel receiver channels. Another possibility for differential operation would be calculation of ToF differences at a non-synchronized Rx if we have two precisely synchronized Tx (time difference of arrival estimation). Finally, we can calculate differences between two synchronized receivers relative to a non-synchronized Tx. In both cases we have to make sure, that at least parts of the distributed Tx or Rx ISAC nodes are absolutely synchronized, e.g. as a part of the infrastructure.
@@
Another coherent receiver operation relates to accurate and high-resolution Doppler estimation. This can again be referred to as a differential operation, as it is based on the relative phase between Tx and Rx. Therefore we need a stable LOS between Tx and Rx, but coherency needs to be maintained only over a limited time interval (also called radar integration time), which maybe as long as one OFDM subframe or even extend over multiple frames. If the ISAC radio nodes are stationary, the estimated target Doppler shift represents the bistatic target speed relative to the pair of nodes. The same requirements applies to background subtraction, SAR imaging and coherent averaging of the received signal. The whole thing gets a little more complicated when the radio nodes are moving since the frequency offset between Tx and Rx LO and the Doppler shift between Tx and Rx cannot be distinguished. However, if the total frequency shift (or phase drift) is compensated, the estimated target-related Doppler corresponds directly to the eDoppler as described above in subsection Multistatic Dynamic Target State Vector Estimation.
@@
The compensation procedure for LOS phase drift requires some comments. The correction should relate exclusively to the LOS path and not be influenced by other multipaths. If we have a sequence of transfer function estimates within the coherent processing interval, we can derive the LOS phase slope along slow time from consecutively estimated CIRs. Within the DFT/FFT framework, the same is achieved from the integral sum of samples of the frequency response function (FRF). However, if there are multiple not resolved dominant reflections, the LOS phase slope will be impaired. In this case, a better estimate can be achieved by at first applying a high-resolution, model-based delay-Doppler estimation approach to the LOS path. Since this implicitly assumes constant speed (a linear phase slope model), some smoothing operation is also included. This may be desirable if the traveling radio nodes perform some wiggling movement that are not relevant for the global trajectory, particularly if it is a drone. Similar is also important for hovering drones, which are assumed to be stationary on average.
VI-B Absolute Synchronization in MS ISAC Networks
@@
At the higher levels of cooperation (IntraSC and InterSC) a common time base and unique time stamps for absolute synchronization is necessary since exchange of processed data takes place. At these levels, position-related measurements such as delay (eToF or bistatic range), Doppler (eDoppler or bistatic velocity), and directional information (DoA, DoD), which are already generated at the lower SLC level, are fused. Therefore, we do no longer have coherency requirements on the delay and carrier phase level. Instead, the remaining requirements arise from multilink access coordination and fusion, target dynamics, and possibly sensor node dynamics.
@@
Precise timing synchronization of distributed clocks is a well known problem in the wide spread field of communication, sensing, and control networks. While GNSS (such as GPS) offers a medium accuracy service (some tens of ns) with global availability, other solutions for more precise synchronization have been proposed based on the IEEE1588 Precision Time Protocol. The White Rabbit (WR) project has demonstrated even sub-ns accuracy synchronization in wired local networks using an Ethernet based protocol [ref39] based on forward-backward transmission.
@@
The infrastructure of 5G networks is already very well synchronized. According to 3GPP/TR 38.816 we can assume a time synchronization of $\pm$ $130\text{\,}\mathrm{ns}\text{/}$ and frequency synchronization of $0.05\text{\,}\mathrm{p}\mathrm{p}\mathrm{m}$, achieved on the basis of GNSS/GPS Primary Reference Time Clocks (PRTC) or Precision Time Protocol, IEEE 1588v2 (PTP). This seems more than enough for data fusion in infrastructure based sensing, where the different sensing sites are stationary (infrastructure only or DL/UL sensing relative to the infrastructure). The situation changes a little when the synchronization master clock of a sensing site is moving, e.g. in case of meshed ICAS established by a moving cloud of drones. The stability of GPS disciplined oscillators (GPSDO) may degrade because of several reasons. Since direct (differential) synchronization at the sensor level is still maintained for every sensing link through the correct use of the reference signal as described above, the time reference for IntraSC data fusion continues to pose no serious problem. While there are a variety of robust GPSDOs available on the market for use in dynamic situations, WR has also proven to be compatible with wireless transmission [ref40]. This would, e.g., allow a sidelink based meshed sensing network to be synchronized relative to the supervising network instance (a remote gNB). Another ad-hoc method for mutual synchronization of distributed radar nodes that does not require a centralized distribution of reference time is described in [ref43]. This all clearly shows that besides of GNSS, we have other approaches at hand that are capable for absolute time synchronization on the fusion level and partly on the PHY-Level. These are applicable also when GNSS is not available or not reliable.
@@
As discussed in the subsection on Multi-Sensor Radio Access and Scheduling we may have broadcast or orthogonal multisensor access. While the broadcast mode allows simultaneous measurement of a subset of distributed MS-MIMO ISAC links (see Figure 10), the orthogonal MS access always deteriorates due to the limitations of the shared channel. FDMA will reduce range resolution, while TDMA influences the availability of measurements in slow time. While the influence on coherent processing (such as ToF or Doppler estimation) has already been discussed, non-isochronous measurements influence the correct assignment of multiple measurement data to the target trajectory. Correct target location requires fusion of isochronous position related data. This may have several issues. Time alignment to achieve coincidence can be achieved by retroactive interpolation or better forward-looking by prediction on the parameter level. Further reasons for data outages here and there can originate from failing target reflections of from target LOS shadowing. In any case, data fusion has to consider timing issues for correct alignment and for labeling age of data, since validity of prediction decreases with time.
@@
VI-C Distributed Data Fusion, Functional Split, Tracking and Sensor Mission Control
@@
Although data fusion and tracking are of paramount importance to the MS ISAC network, we can only provide very limited, general insights here. The aim of data fusion is to combine related measurements to infer about the dynamic target state vector. Tracking takes into account temporal evolution of measurements and finishes estimation by smoothing and error correction. The underlying network architecture suggests a multilevel approach for both fusion and tracking. Although one of the main advantages of ISAC is that it provides a network for transport and computing resources for data fusion, this raises several questions. The transport links may have different capacity and latency performance depending on the available medium, which can reach from a mobile or fixed wireless link to an optical RF over Fiber (RFoF) link with highest performance. There can also be significant differences in terms of computing resources, ranging from low performance in small drones over medium performance in cars up to high performance in base stations. The Open RAN paradigm seems promising for ISAC data transport and control flow routing within an infrastructure site due to its transparent and multi-vendor approach, which allows interoperability of components at different levels, also including specialized sensing hardware such as dedicated SDR based sniffers or high performance computing instances. Sensing functional split should consider availability and allocation of resources for both, communication and computation, see [ref41, ref66].
@@
The obvious, albeit rather conventional approach, is to detect the targets echos on the sensor level and transmit the preprocessed data set to the fusion center where further fusion takes place. This effectively establishes a multilayer approach for data fusion. However, the hierarchy of fusion needs to be defined. For example, it may happen that we have incomplete measurements at the lower level that nevertheless can contribute to the final result. Another problem is data association, which means that we may not know (especially at the lower level) which measurements belong to the same target. Eventually, the capacity and latency of the data links between the sensors and sensor clusters and the fusion may vary, which would have a significant impact on the functional split of data fusion across the different levels of the ISAC network. In distributed MS ISAC there are specific reasons to decide about the functional split of fusion on the sensor level (SLC) or on network level (IntraSC or InterSC). Local fusion on the sensor level enables early evaluation of the sensor performance, thereby supporting decisions regarding the relevance of the respective node and its data for the final fusion. An overdetermined set of measurements (excessive number of links) can reduce variance, increase robustness to missing information and help to sort out unreliable and probably wrong measurements. In this way, the network can be relieved of the transmission and processing of irrelevant data to the fusion center. Data reduction can have decisive influence in cases of wireless transport links with their varying capacity and latency. In addition to the impact to location accuracy (GDOP), the selection of sensors for fusion would also depend on the sensor configuration such as available bandwidth, frequency, power, antenna and array configuration, etc. The immediate evaluation of local data can help in deciding on the allocation of local radio resources and on the sensor mission. The latter seems to be most important in cases of movable, mobile or otherwise reconfigurable ISAC radio nodes to decide about the future sensor mission. Moreover, control about local information to be shared can also support keeping privacy and technical integrity of sensing entities.
@@
It is interesting to see that AI based data fusion approaches also make a general distinction between early and late fusion, see [ref57]. While early fusion, which is primarily sought after in deep learning, attempts to combine data at a raw data input level, late fusion begins combining data after independent processing on the sensor level. Early fusion is often considered to be more powerful. On the other hand, late fusion preserves more modularity and structural flexibility. Deeper research into communication and computation resources constraints for early and late AI based data fusion would be interesting.
@@
Tracking accounts for the movement of targets relative to senor nodes. It establishes the temporal and spatial connection of hypothesized parameters. Tracking takes advantage of target inertia for smoothing and prediction of noisy and otherwise erroneous estimates. The Kalman paradigm [bar2004estimation] recursively updates the target state against disturbed observations. The dynamic state vector can include a more or less constrained model in terms of 3D position, orientation, speed, acceleration, etc. Target tracking helps to sort out wrong position hypotheses and infer about correct data association in multi-target tracking by Multi-Hypotheses Tracker (MHT), [reid1979MultiHypothesisTracker] or Joint Probabilistic Data Association (JPDA), [barshalom2009JPDAMultiTargetTrackin] filtering. Since tracking can already take place on the lower level with respect to position related parameters like ToF of Doppler shift, it can help to pre-process data on the sensor level, contribute to data reduction and help to predictively allocate resources on the sensor level.
@@
Fusion and tracking on the intersite level aims more on a global picture in spacious area which is covered by multiple independent (perhaps not overlapping) sensor sites. Cooperation at this level no longer involves direct sensor data fusion. Instead, it involves the fusion of multiple local scenarios of dynamic target state vectors and trajectories. Global perception and situation recognition are generated and made available to users, or even used by the ISAC network for cognitive conclusions about actions to be taken. The latter also opens doors for cognitive ISAC. This means that sensing targets are identified at the global scenario level, translated into local sensing missions, and then sent as requests to the lower levels of the multi-layer ISAC network, where decisions about the allocation of resources are made. Therefore, crosslayer signaling and cooperation between the sensing layers (rsp. levels) are important.

### v2

@@ VI Synchronization, Data Fusion, and Tracking in Distributed MS ICAS Networks
Advanced MS-ISAC is highly cooperative. This begins with the bistatic sensing architecture, which is quite close to the basic radio access scheme of mobile communication. It already has all the advantages of bistatic radar with respect to arget detection and identification. We refer to this first level as (i) Sensor Level Cooperation (SLC), which refers to processes that can be performed by each sensor pair independently of each other. Although it does not involve the exchange of processed data with other sensor nodes, it is already cooperative, e.g., when the sensor receives measurement signals from several illuminating radio nodes, e.g., in the MS-ISAC up-link mode. The next level of cooperation is achieved when multiple sensing links share processed data to estimate and track the 3D dynamic target state vector of one or more targets. This level already provides important benefits of MS-ISAC, which include spatially smooth performance distribution by multisensor data fusion, multilink access coordination, and local radio resource allocation. According to our initial discussion on MS-ISAC architectures, the ISAC nodes may belong to a distributed infrastructure as an MS sensing cluster site, include mobile UEs (in DL or UL), or belong to an independent, self-contained cloud of meshed radio nodes. We refer to this level as (ii) Intra Site (or Intra Cloud) Cooperation (IntraSC). The third level of cooperation is called (iii) Intersite (or Inter Cloud) Cooperation (InterSC). It fuses the estimated target state vectors from different sensing sites or sensing clouds and can combine or fuse ISAC data with other (including non ISAC) information to produce a comprehensive situation report. In the sequel, we will at first discuss some synchronization issues and then shortly refer to distributed data fusion and tracking in MS-ISAC.
VI-A Differential Coherent Operation in MS-ISAC at Sensor Level
@@
At the sensor level, parameters are estimated that are closely related to the dynamic target state vector, such as time delay (rsp. range), Doppler frequency shift (rsp. relative speed), or directions (DoD, DoA). While the latter in general requires more complicated RF interfaces (because of the antenna arrays), the former two require synchronization with respect to sampling and Local Oscillator (LO) frequencies offsets of the bistatic radio nodes [ref90]. Array processing is easier in this respect as only local (or platform-related) coherent processing is necessary. Therefore, distributed MS-MIMO ISAC is often claimed to have disadvantages due to a lack of synchronization between sensor nodes. On closer inspection, however, many solutions can be found if we avoid excessive and unnecessary requirements. As already explained above, MS-MIMO ISAC requires special efforts to establish sensor access coordination and synchronization. However, since the Tx and Rx are connected as a communication transceiver pair in a multiuser or mesh communication network, the receiving radio nodes are already basically synchronized to the transmitter by using the Primary and Secondary Synchronization Signals (PSS, SSS) [ref91, ref92]. Therefore, multisensor operation, OFDM frame coordination, and carrier frequency adjustment are already accomplished according to the requirements for communication. This ensures orthogonal, mutual interference-free medium access and synchronous FFT processing, which is a fundamental advantage for low variance system identification [ref30]. This is further supported by the design of the OFDM symbol, which is not chosen too long to avoid time selectivity and supplemented by a cyclic prefix. For further synchronization update, one to four Demodulation and Phase Tracking Reference Signals (PT-RS and DM-RS) can be allocated per slot within the Physical Downlink / Uplink Shared Channel (PDSCH / PUSCH) [ref91, ref92].
@@
Although these timing and clock synchronization schemes are very flexible and powerful for communication even in situations with high Doppler shift, they seem not to be enough for ISAC on the sensor level cooperation. ToF and Doppler estimation call for sub-ns delay and precise carrier phase synchronization (the latter becoming more demanding the higher the frequency). Fortunately, these problems can be mitigated by carrying out differential measurements. A striking historical example is given by the well-known passive radar concept, which uses a “transmitter of opportunity” as an illuminator and a “passive” (receiving only) sensor. PCL does not presume any explicit signaling scheme for synchronization between Tx and Rx. The bistatic eToF is estimated from the received communication signal, which is correlated against the LOS-reference. Classical passive radar receivers use two synchronous channels, an omnidirectional surveillance channel and a directive channel receiving the clean LOS reference, which must be free of multipath. In case of the more advanced CPCL, the eToF and eDoppler can be actually calculated from the same CIR by the delay and frequency difference between LOS and BTP, provided that the correlation reference is correctly reconstructed and LOS and target peaks are clearly identified. This eliminates the remaining Tx and Rx clock and LO offset and even avoids drift and deviations between parallel receiver channels. Another possibility for differential operation would be the calculation of ToF differences at a non-synchronized Rx if we have two precisely synchronized Tx (time difference of arrival estimation). Finally, we can calculate differences between two synchronized receivers relative to a non-synchronized Tx. In both cases, we have to make sure that at least parts of the distributed Tx or Rx ISAC nodes are absolutely synchronized, e.g., as a part of the infrastructure.
@@
Another coherent receiver operation relates to accurate and high-resolution Doppler estimation. This can again be referred to as a differential operation, as it is based on the relative phase between Tx and Rx. Therefore, we need a stable LOS between Tx and Rx, but coherency needs to be maintained only over a limited time interval (also called radar integration time), which may be as long as one OFDM subframe or extend over multiple frames. If the ISAC radio nodes are stationary, the estimated target Doppler shift represents the bistatic target speed relative to the pair of nodes. The same requirements apply to background subtraction, Synthetic Aperture Radar (SAR) imaging, and coherent averaging of the received signal. The whole thing gets a little more complicated when the radio nodes are moving, since the frequency offset between Tx and Rx LO and the Doppler shift between Tx and Rx cannot be distinguished. However, if the total frequency shift (or phase drift) for the composite sum of all received paths is compensated, the estimated target-related Doppler corresponds directly to the eDoppler as described above in subsection Multistatic Dynamic Target State Vector Estimation.
@@
The compensation procedure for LOS phase drift requires some comments. The correction should relate exclusively to the LOS path and not be influenced by other multipaths. If we have a sequence of transfer function estimates within the coherent processing interval, we can derive the LOS phase slope along slow time from consecutively estimated CIRs. Within the Discrete Fourier Transform (DFT)  / FFT framework, the same is achieved from the integral sum of samples of the frequency response function (FRF). However, if there are multiple unresolved dominant reflections, the LOS phase slope will be impaired. In this case, a better estimate can be achieved by first applying a high-resolution, model-based delay-Doppler estimation approach to the LOS path. Since this implicitly assumes constant speed (a linear phase slope model), some smoothing operation is also included. This may be desirable if the traveling radio nodes perform some wiggling movements that are not relevant for the global trajectory, particularly if it is a drone. Similar is also important for hovering drones, which are assumed to be stationary on average.
VI-B Absolute Synchronization in MS-ISAC Networks
@@
At the higher levels of cooperation (may it be IntraSC or InterSC), a common time base and unique time stamps for absolute synchronization are necessary since the exchange of processed data takes place. At these levels, position-related measurements such as delay (eToF or bistatic range), Doppler (eDoppler or bistatic velocity), and directional information (DoA, DoD), which are already generated at the lower SLC level, are fused. Therefore, we no longer have coherency requirements on the delay and carrier phase level. Instead, the remaining requirements arise from multilink access coordination and fusion, target dynamics, and possibly sensor node dynamics which is on a relaxed scale.
@@
Precise timing synchronization of distributed clocks is a well-known problem in the widespread field of communication, sensing, and control networks. While GNSS (such as GPS) offers a medium accuracy service (some tens of ns) with global availability, other solutions for more precise synchronization have been proposed based on the IEEE1588 Precision Time Protocol. The White Rabbit (WR) project has demonstrated even sub-ns accuracy synchronization in wired local networks using an Ethernet-based protocol [ref39] based on forward-backward transmission.
@@
The infrastructure of 5G networks is already very well synchronized. According to 3GPP/TR 38.816, we can assume a time synchronization of $\pm$ $130\text{\,}\mathrm{ns}\text{/}$ and frequency synchronization of $0.05\text{\,}\mathrm{p}\mathrm{p}\mathrm{m}$, achieved based on GNSS/GPS Primary Reference Time Clocks (PRTC) or Precision Time Protocol, IEEE 1588v2 (PTP). This seems more than enough for data fusion in infrastructure-based sensing, where the different sensing sites are stationary (infrastructure only or DL  / UL sensing relative to the infrastructure). The situation changes a little when the synchronization master clock of a sensing site is moving, e.g., in the case of a meshed ISAC established by a moving cloud of drones. The stability of GPS disciplined oscillators (GPSDO) may degrade for several reasons. Since direct (differential) synchronization at the sensor level is still maintained for every sensing link through the correct use of the reference signal as described above, the time reference for IntraSC data fusion continues to pose no serious problem. While there are a variety of robust GPSDOs available on the market for use in dynamic situations, WR has also proven to be compatible with wireless transmission [ref40]. This would, e.g., allow a sidelink-based meshed sensing network to be synchronized relative to the supervising network instance (a remote gNB). Another ad-hoc method for mutual synchronization of distributed radar nodes that does not require a centralized distribution of reference time is described in [ref43]. This all clearly shows that besides GNSS, we have other approaches at hand that are capable of absolute time synchronization on the fusion level and partly on the PHY-Level. These are also applicable when GNSS is not available or not reliable.
@@
As discussed in the subsection on Multi-Sensor Radio Access and Scheduling, we may have broadcast or orthogonal multisensor access. While the broadcast mode allows simultaneous measurement of a subset of distributed MS-MIMO ISAC links (see Figure 10), the orthogonal MS access always deteriorates due to the limitations of the shared channel. FDMA will reduce range resolution, while TDMA influences the availability of measurements in slow time. While the influence on coherent processing (such as ToF or Doppler estimation) has already been discussed, non-isochronous measurements influence the correct assignment of multiple measurement data to the target trajectory. Correct target location requires fusion of isochronous position-related data. This may have several issues. Time alignment to achieve coincidence can be achieved by retroactive interpolation or, better, forward-looking by prediction on the parameter level. Further reasons for data outages here and there can originate from failing target reflections or from target LOS shadowing. In any case, data fusion has to consider timing issues for correct alignment and for labeling the age of data, since the validity of the prediction decreases with time.
@@
VI-C Distributed Data Fusion, Functional Split, Tracking, and Sensor Mission Control
@@
Although data fusion and tracking are of paramount importance to the MS-ISAC network, we can only provide very limited, general insights here. Data fusion aims to combine related measurements to infer about the dynamic target state vector. Tracking takes into account the temporal evolution of measurements and finishes estimation by smoothing and error correction. The underlying network architecture suggests a multilevel approach for both fusion and tracking. Although one of the main advantages of ISAC is that it provides a network for transport and computing resources for data fusion, this raises several questions. The transport links may have different capacity and latency performance depending on the available medium, which can range from a mobile or fixed wireless link to an optical RF over Fiber (RFoF) link with the highest performance. There can also be significant differences in terms of computing resources, ranging from low performance in small drones to medium performance in cars up to high performance in base stations. The Open RAN paradigm seems promising for ISAC data transport and control flow routing within an infrastructure site due to its transparent and multi-vendor approach, which allows interoperability of components at different levels, also including specialized sensing hardware such as dedicated SDR-based sniffers or high-performance computing instances. Sensing functional split should consider availability and allocation of resources for both communication and computation, see [ref41, ref66].
@@
The obvious, albeit rather conventional approach, is to detect the targets echos on the sensor level and transmit the preprocessed data set to the fusion center, where further fusion takes place. This effectively establishes a multilayer approach for data fusion. However, the hierarchy of fusion needs to be defined. For example, we may have incomplete measurements at the lower level that nevertheless can contribute to the final result. Another problem is data association, which means that we may not know (especially at the lower level) which measurements belong to the same target. Eventually, the capacity and latency of the data links between the sensors and sensor clusters and the fusion may vary, which would have a significant impact on the functional split of data fusion across the different levels of the ISAC network. In distributed MS-ISAC there are specific reasons to decide about the functional split of fusion on the sensor level (SLC) or on the network level (IntraSC or InterSC). Local fusion on the sensor level enables early evaluation of the sensor performance, thereby supporting decisions regarding the relevance of the respective node and its data for the final fusion. An overdetermined set of measurements (excessive number of links) can reduce variance, increase robustness to missing information, and help to sort out unreliable and probably wrong measurements. In this way, the network can be relieved of the transmission and processing of irrelevant data to the fusion center. Data reduction can have a decisive influence in cases of wireless transport links with their varying capacity and latency. In addition to the impact on location accuracy (GDoP), the selection of sensors for fusion would also depend on the sensor configuration, such as available bandwidth, frequency, power, antenna, and array configuration, etc. The immediate evaluation of local data can help in deciding on the allocation of local radio resources and on the sensor mission. The latter seems to be most important in cases of movable, mobile, or otherwise reconfigurable ISAC radio nodes to decide about the future sensor mission. Moreover, control over local information to be shared can also support keeping privacy and the technical integrity of sensing entities.
@@
It is interesting to see that AI-based data fusion approaches also make a general distinction between early and late fusion, see [ref57]. While early fusion, which is primarily sought after in deep learning, attempts to combine data at a raw data input level, late fusion begins combining data after independent processing on the sensor level. Early fusion is often considered to be more powerful. On the other hand, late fusion preserves more modularity and structural flexibility. Deeper research into communication and computation resource constraints for early and late AI-based data fusion would be interesting.
@@
Tracking accounts for the movement of targets relative to sensor nodes. It establishes the temporal and spatial connection of hypothesized parameters. Tracking takes advantage of target inertia for smoothing and prediction of noisy and otherwise erroneous estimates. The Kalman paradigm [bar2004estimation] recursively updates the target state against disturbed observations. The dynamic state vector can include a more or less constrained model in terms of 3D position, orientation, speed, acceleration, etc. Target tracking helps to sort out wrong position hypotheses and infer about correct data association in multi-target tracking by Multi-Hypotheses Tracker (MHT) [reid1979MultiHypothesisTracker] or Joint Probabilistic Data Association (JPDA) filtering  [barshalom2009JPDAMultiTargetTrackin]. Since tracking can already take place on the lower level with respect to position-related parameters like ToF of Doppler shift, it can help to pre-process data on the sensor level, contribute to data reduction, and help to predictively allocate resources on the sensor level.
@@
Fusion and tracking on the intersite level aims more at a global picture in a spacious area which is covered by multiple independent (perhaps not overlapping) sensor sites. Cooperation at this level no longer involves direct sensor data fusion. Instead, it involves the fusion of multiple local scenarios of dynamic target state vectors and trajectories. Global perception and situation recognition are generated and made available to users, or even used by the ISAC network for cognitive conclusions about actions to be taken. The latter also opens doors for cognitive ISAC. This means that sensing targets are identified at the global scenario level, translated into local sensing missions, and then sent as requests to the lower levels of the multi-layer ISAC network, where decisions about the allocation of resources are made. Therefore, cross-layer signaling and cooperation between the sensing layers (rsp. levels) are important.
The hierarchical concept of data fusion also supports division of work and interoperability between different operators if infrastructure related sensing is concerned (infrastructure-only or UL/DL sensing). Since a sensor cluster usually belongs to a base station, it is operated by a mobile network operator. This includes sensor configuration, maintenance, performance management, and data compression. Sensing results with well-defined quality can be further delivered via a customized interface to a higher level fusion center, e.g. to aggregate nation-wide situation maps or to monitor extended locations of critical infrastructure for safety critical application.